A refrigerator has a coefficient of performance equal to 5.8. If the refrigerator absorbs 262.9 J of thermal energy from a cold reservoir in each cycle, find the work done on the system each cycle (in J). Enter a number with one digit behind the decimal point.
Answer #1

Qin = 262.9J and Cop = 5.8

Since Cop = Qin/W

=) W = Qin/Cop = 262.9/5.8 = 45.3 Joule.

That means work done per cycle W = 45.3 J